# Imitative Membership Inference Attack

### v1

@@
A Membership Inference Attack (MIA) assesses how much a target machine learning model reveals about its training data by determining whether specific query instances were part of the training set. State-of-the-art MIAs rely on training hundreds of shadow models that are independent of the target model, leading to significant computational overhead. In this paper, we introduce Imitative Membership Inference Attack ($\mathsf{IMIA}$), which employs a novel imitative training technique to strategically construct a small number of target-informed imitative models that closely replicate the target model’s behavior for inference. Extensive experimental results demonstrate that $\mathsf{IMIA}$ substantially outperforms existing MIAs in various attack settings while only requiring less than 5% of the computational cost of state-of-the-art approaches [^1] Our code is available at https://github.com/zealscott/IMIA..
@@ 1 Introduction
Over the past decade, machine learning (ML) has seen remarkable advances, with models such as neural networks increasingly being trained on sensitive datasets. This trend has raised growing concerns about the privacy risks associated with these models. Membership inference attacks (MIAs) [48] have been proposed to measure the degree to which a model leaks information by determining whether some instances were part of its training set. Closely related to Differential Privacy (DP) [13, 32], MIAs have become a widely adopted technique for empirical auditing of various trustworthy risks in ML models [39, 51, 56, 57, 42], and serve as key components for more sophisticated attacks [4, 5, 52].
MIAs in the black-box setting [59, 50, 58, 3] typically exploit the model’s behavioral discrepancy between its training instances (i.e., members) and non-training instances (i.e., non-members) for inference. A widely-used strategy to explore the discrepancy is shadow training [48], which involves training multiple shadow models on datasets drawn from the same distribution as the target model’s training set. For each query instance, the membership scores generated by the shadow models trained without the instance (shadow out models) represent the out distribution for the instance. In some settings, one also has shadow models that are trained with the instance as a member, and can use the scores generated by these shadow in models to compute an in distribution. Given a target model, one tries to tell which distribution the score from the target model fits better. MIAs [54, 19, 9, 60, 3] utilizing shadow training have shown strong attack performance.
@@
A critical limitation of existing shadow-based MIAs lies in their substantial computational overhead. Training each shadow model incurs non-trivial overhead, and state-of-the-art attacks like LiRA [3] and PMIA [9] require training hundreds of shadow models to estimate the likelihood ratio for inference. This requirement imposes a substantial computational burden, which reduces the feasibility of using these state-of-the-art MIAs for practical privacy auditing [60]. In addition, it also impedes research reproducibility within the privacy community (see Section˜2 for further discussion).
We observe that this inefficiency stems from the target-agnostic design of shadow training. That is, the current shadow training process fails to take advantage of knowing the target model under attack. As a result, the shadow models learn only the general patterns of members and non-members, rather than the specific behavioral discrepancies of the target model. As shown in the top row of Figure˜1, this target-agnostic approach causes shadow models to exhibit high predictive variance across instances with varying levels of attack vulnerability. Consequently, existing MIAs need to train a multitude of shadow models to capture this variability, resulting in both substantial computational overhead and suboptimal performance.
@@
We note that some attacks [58, 15, 30] employ model distillation to construct shadow models from the target model. However, these methods directly apply the standard distillation technique [20] without tailoring it to capture the membership-related behaviors of the target model. Furthermore, the resulting distillation models can only reflect behaviors on non-member instances, making it infeasible to exploit behavioral discrepancies for effective membership inference. A more detailed discussion of the differences between our approach and distillation-based methods is provided in Section˜3.2.
We compare the proposed attack with a broad range of MIAs on six benchmark datasets (4 image and 2 non-image datasets), under both non-adaptive and adaptive settings [9] (also referred to as “offline” and “online” [3]; see Section˜2 for details). Experimental results show that $\mathsf{IMIA}$ consistently outperforms all evaluated attacks (including distillation-based attacks) while using significantly fewer shadow models, with particularly strong gains in the increasingly advocated non-adaptive setting. For example, $\mathsf{IMIA}$ surpasses the state-of-the-art non-adaptive attack (i.e., PMIA [9]) by over $\mathbf{9\times}$ in true positive rate at a 0.001% false positive rate on Fashion-MNIST, while using only 5% of its computational resources. Similarly, $\mathsf{IMIA}$ outperforms all adaptive attacks with far less computation. We also conduct comprehensive analyses to show the effectiveness of imitative training over shadow training, evaluate the impact of components in $\mathsf{IMIA}$, and assess the effectiveness of existing defenses against the proposed attack. In summary, we make the following contributions:
@@ 2 Problem Definition and Threat Models
The goal of a membership inference attack (MIA) [48] is to determine whether some instances were included in the training data of a given model $f_{\theta}$. In this paper, we consider the model to be a neural network classifier $f_{\theta}:\mathcal{X}\to\Delta^{c}$, where $f_{\theta}$ is a learned function that maps an input data sample $x\in\mathcal{X}$ to a probability distribution over $c$ classes, where $\Delta^{c}$ denotes the $c$-dimensional simplex. Given a dataset $D$, we denote by $f_{\theta}\leftarrow\mathcal{T}(D)$ the process of training a model $f_{\theta}$, parameterized by weights $\theta$, using a training algorithm $\mathcal{T}$ on dataset $D$.
In early literature [59, 24], membership inference is defined via a security game in which the adversary is asked to determine the membership of a single instance. Recent work [9] refined this definition to more precisely characterize the adversary’s capabilities, particularly whether the adversary is allowed to train shadow models using query instances. This refinement enables a clear distinction between non-adaptive (i.e., offline) and adaptive (i.e., online) settings [3], which often involve different attack strategies and yield significantly different attack performance. Thus, we follow [9] and define membership inference through the following security game:
Definition 1 (Membership Inference Security Game [9]).
@@
    The challenger randomly selects a subset $D_{\mathrm{a}}\subseteq D$ and samples a set $D_{\mathrm{b}}$ from $\mathbb{D}$. These two sets are combined to create a query set: $D_{\mathrm{query}}=D_{\mathrm{a}}\cup D_{\mathrm{b}}$, which the challenger then sends to the adversary.
@@
In this paper, we focus on membership inference attacks in black-box scenarios, where the adversary is granted oracle access to the target model $f_{\theta}$, but is not given its parameters. Specifically, the adversary can query the model to obtain softmax output probabilities (using the default softmax temperature of 1) for any input instance. State-of-the-art MIAs [3, 54, 19, 60, 9] leverage shadow models [48] to analyze how the model’s outputs change in response to discrepancies between its members versus non-members. We assume that the adversary can construct a dataset by sampling from the data distribution $\mathbb{D}$, and use subsets of this dataset to train shadow models. Depending on when the shadow models are trained, we consider the following two threat models.
@@
Non-Adaptive Setting. In this setting, the adversary is only allowed to train shadow models before the adversary learns the query set $D_{\text{query}}$ (i.e., before step 2 in Definition˜1). That is, the adversary constructs an attack dataset $D_{\text{adv}}^{\text{non-adapt}}$ (from which the adversary will sample subsets to train shadow models) by sampling from $\mathbb{D}$, i.e., $D_{\text{adv}}^{\text{non-adapt}}\sim\mathbb{D}$. For most practical classification tasks, the sizes of $D_{\text{adv}}^{\text{non-adapt}}$ and $D_{\text{query}}$ are relatively small compared to the entire data distribution $\mathbb{D}$, meaning the probability of each instance belonging to both datasets is quite low. As a result, the adversary can only observe the model’s behavior when a query instance is not part of the training set, leading to shadow out models available for membership inference.
@@
This non-adaptive (offline) setting has attracted growing interest from recent studies [2, 30, 60, 19] because it models a realistic attack scenario: the adversary receives a sequence of membership queries and must respond without incurring the cost of retraining shadow models for each query. In this paper, we also follow these studies and mainly focus on the non-adaptive setting.
@@
Adaptive Setting. In contrast, under the adaptive setting, the adversary is allowed to train shadow models after receiving the query set $D_{\text{query}}$ (i.e., after step 2 in Definition˜1). Thus, the adversary samples $D^{\prime}$ from $\mathbb{D}$ and combines it with $D_{\text{query}}$ to create a dataset for shadow training, i.e., $D_{\text{adv}}^{\text{adapt}}\leftarrow D^{\prime}\cup D_{\text{query}}$. This allows the adversary to train shadow in models (trained with the instance) and shadow out models (trained without the instance) for each query instance, and exploit their behavioral differences for membership inference.
@@
Recent studies [60, 2, 43] criticize this setting as unrealistic in practice, as it requires training new shadow models for each batch of query instances. Nevertheless, we demonstrate in Section˜3.1 that our proposed method can be adapted to the adaptive setting, achieving state-of-the-art performance while significantly reducing computational overhead.
@@
High Computational Cost in Both Settings. State-of-the-art MIAs in both settings require training a large number of shadow models to achieve their best performance. For instance, LiRA [3] in the adaptive setting and PMIA [9] in the non-adaptive setting both suggest training 256 shadow models. The practical implication of this is severe: a single LiRA run on CIFAR-10 can take around 6 days on a modern A100 GPU using its official implementation. Such high computational demands make it infeasible for practical privacy auditing [60] and difficult to reproduce research results. It has become common practice for subsequent studies [30, 54, 19] to use a computationally cheaper version of LiRA (e.g., with 64 or 128 shadow models) as their strongest baseline for evaluation, acknowledging the original version is too costly to run. However, since LiRA’s performance degrades significantly with fewer shadow models (as shown in [60] and evidenced by our experiments), this compromise leads to skewed evaluations against a weakened baseline.
Missed Opportunities of Existing Attacks. The adversary in MIAs has query access to the target model, but state-of-the-art MIAs [54, 19, 60, 3] rely on a target-agnostic shadow training process [48]. This means shadow models are trained independently of the target model, learning only general patterns of members and non-members across models. We argue for a target-informed approach, where shadow models should incorporate knowledge of the target model to better capture its behavioral discrepancies for more effective attack.
@@ 3 Imitative Membership Inference Attack
Motivation of Imitative MIA. Our work is motivated by a core limitation in most MIAs: the high predictive variance of their target-agnostic shadow models. Formally, let $z(x)\leftarrow g(x,f_{\psi})$ denote the membership signals that the adversary exploits from a shadow model $f_{\psi}$ to infer the membership of an instance $x$. Let $o_{\theta}\leftarrow\mathcal{Q}(f_{\theta})$ be the information the adversary can observe from the target model $f_{\theta}$, which consists of the target’s output confidence vectors (softmax output probabilities) on available data instances. According to the Law of Total Variance [44], conditioning on the target’s outputs reduces the variance of membership signals:
$\mathbb{E}_{o_{\theta}}\left[\operatorname{Var}\left(z(x)\mid o_{\theta}\right)\right]\leq\operatorname{Var}[z(x)].$
This inequality indicates that, in an average sense, membership signals become more stable when informed by the target model’s outputs; and accordingly, a higher adversarial success rate is expected when their strategy is properly conditional on the target model. Motivated by this insight, we propose a new shadow training technique, imitative training, which leverages the target model’s knowledge to train more informative shadow models (i.e., imitative models) that explicitly mimic the target’s behavior. This design yields more informative membership signals and, ultimately, more effective attacks. The next subsection describes our approach in detail, and Section˜4.3 provides empirical evidence of its effectiveness.
@@ 3.1 Attack Method
Attack Overview. At its core, $\mathsf{IMIA}$ utilizes a novel imitative training approach to train target-informed imitative models that imitate the behavior of the target model. We begin by introducing the membership signal and the loss function used for imitative training. Next, we outline the imitative training process. Finally, we present the complete workflow of $\mathsf{IMIA}$, and detail how it can be applied to the adaptive setting.
Scaled Confidence Score. We follow [3] and define the membership signal of query instance $(x,y)$ on model $f$ as:
@@
where $f(x)_{y}$ denotes the model’s output probability for the true class $y$, and $\max_{y^{\prime}\neq y}f(x)_{y^{\prime}}$ is the highest probability among all incorrect classes. This signal captures the model’s prediction confidence and is widely used in prior MIAs [3, 54, 9], as it is a more effective indicator of membership than alternatives such as output loss.
@@
0: Target model $f_{\theta}$, imitation dataset $D_{\text{imitate}}$, pivot dataset $D_{\text{pivot}}$, epochs $T_{1},T_{2}$, learning rate $\eta$
@@
5:   $L_{\text{imitate}}\leftarrow\frac{1}{|B|}\sum_{(x,y)\in B}\mathcal{L}_{\text{imitate}}(x,y,f_{\psi},f_{\theta})$ $\triangleright$ Eq. (2)
6:   $\psi\leftarrow\psi-\eta\nabla_{\psi}L_{\text{imitate}}$
7:  end for
8:end for
[…]
10:# continue training to obtain imitative in model for pivots
11:for epoch $1,\dots,T_{2}$ do
12:  for each batch $B$ in $D_{\text{pivot}}$ do
13:   $L_{\text{ce}}\leftarrow\frac{1}{|B|}\sum_{(x,y)\in B}\mathcal{L}_{\text{ce}}(x,y,f_{\psi})$ $\triangleright$ cross-entropy loss
14:   $\psi\leftarrow\psi-\eta\nabla_{\psi}L_{\text{ce}}$
15:  end for
@@
18:return  $f_{\text{out}},\ f_{\text{in}}^{\text{pivot}}$
@@
Here, the weighting scheme prioritizes the probability corresponding to the ground-truth class, $y$, and the most likely incorrect class, as predicted by the target model. This design encourages imitative models to focus on the most indicative signals for membership, which aligns with our attack signal described in Equation˜1.
Our approach can be interpreted as a weighted version of logits matching in model distillation [25], with a weighting strategy specifically tailored to membership inference. In Section˜4.4, we compare it to the commonly used Kullback-Leibler (KL) divergence [20] and also explore other formulations of the imitation loss to demonstrate its effectiveness. A more detailed discussion of the connections to model distillation is provided in the next subsection.
Imitative Training. The imitative training procedure is outlined in Algorithm˜1. It requires query access to the target model and utilizes two datasets selected from the adversary’s data to mimic the target model’s behavior: an imitation dataset $D_{\text{imitate}}$ and a pivot dataset $D_{\text{pivot}}$ (to be detailed later). The training proceeds in two phases. First, we use the proposed imitation loss to train a model for $T_{1}$ epochs to match the target model’s outputs on $D_{\text{imitate}}$ (lines 3–9). Since query instances are unlikely to appear in the adversary’s dataset in the non-adaptive setting, the resulting model serves as an imitative out model, imitating how the target model behaves on non-member inputs. We then continue training the model for $T_{2}$ epochs on a pre-selected pivot dataset using the standard cross-entropy loss (lines 11–16). This yields an imitative in model for the pivot data, whose behaviors on these instances will serve as proxies to approximate how the target model behaves when a query instance is in the training set.
@@
Our computation of the score $\Lambda$ can be interpreted as a Gaussian likelihood ratio test [3] with fixed variance, which allows us to frame the inference task as a standard hypothesis testing for membership inference. Note that the prepare phase can be done before accessing query instances and only needs to be performed once to answer any membership queries using the same set of imitative models. During the inference phase, no additional models are trained. The efficiency analysis of each phase is detailed in Section˜4.2.
Selecting Pivot Data and Finding Proxies. Our attack requires a pivot dataset to train imitative in models. This dataset is also used to select proxy instances to approximate the behavior of a query instance belonging to the target model’s training set. An effective pivot dataset should be diverse to provide representative proxies for a wide range of membership queries and enable fast training of imitative models. To construct such a dataset, we adopt a simple yet effective selection strategy. We query all instances in $D_{\text{adv}}^{\text{non-adapt}}$ on the target model $f_{\theta}$ and compute their prediction losses. Then, from each class, we select $k$ instances with the lowest loss. These selected samples collectively form $D_{\text{pivot}}$. We set $k=100$ by default, and our ablation study in Section˜4.4 shows that our attack is robust to this choice. During inference, we follow [9] and select proxies for a query instance by retrieving all samples from $D_{\text{pivot}}$ that belong to the same class.
@@
$\mathsf{IMIA}$ in the Adaptive Setting. While $\mathsf{IMIA}$ is primarily designed for the non-adaptive setting, it can also be applied to the adaptive attack setting. As shown in Algorithm˜3, for a query instance $(x,y)$, we train $N$ pairs of imitative in models (trained with $(x,y)$) and imitative out models (trained without $(x,y)$) on random subsets of the adversary’s dataset $D_{\text{adv}}^{\text{adapt}}$ (lines 2–9). This procedure directly learns the target model’s behavior when the query instance is in its training set, thus eliminating the need for a pivot dataset and proxy selection. The final score is computed by comparing the query’s confidence score on the target model to the mean scores of the imitative in and out distributions (lines 11-14). In addition, while the algorithm is described for a single query, the procedure can be easily parallelized across all query instances using the same set of imitative models. Specifically, we follow [3] and construct the subsets such that each instance $(x,y)\in D_{\text{query}}$ appears in $N/2$ subsets, and train $N$ imitative models on these subsets. This ensures that the same $N$ imitative models are used to infer the membership for all instances in $D_{\text{query}}$.

### v2

@@
A Membership Inference Attack (MIA) assesses how much a target machine learning model reveals about its training data by determining whether specific query instances were part of the training set. State-of-the-art MIAs rely on training hundreds of shadow models that are independent of the target model, leading to significant computational overhead. In this paper, we introduce Imitative Membership Inference Attack ($\mathsf{IMIA}$), which employs a novel imitative training technique to strategically construct a small number of target-informed imitative models that closely replicate the target model’s behavior for inference. Extensive experimental results demonstrate that $\mathsf{IMIA}$ substantially outperforms existing MIAs in various attack settings while only requiring less than 5% of the computational cost of state-of-the-art approaches.
@@ 1 Introduction
Over the past decade, machine learning (ML) has seen remarkable advances, with models such as neural networks increasingly being trained on sensitive datasets. This trend has raised growing concerns about the privacy risks associated with these models. Membership inference attacks (MIAs) [sp17miashokri] have been proposed to measure the degree to which a model leaks information by determining whether some instances were part of its training set. Closely related to Differential Privacy (DP) [dwork_dp, ccs13membership], MIAs have become a widely adopted technique for empirical auditing of various trustworthy risks in ML models [arxiv20privacy_meter, tensorflow_mia, dsi, memorization, copyright], and serve as key components for more sophisticated attacks [usenix21llm_extract, usenix23diffusion_extract, iclr25usage].
MIAs in the black-box setting [csf18privacy, usenix21systematic, ccs22enhanced, sp22lira] typically exploit the model’s behavioral discrepancy between its training instances (i.e., members) and non-training instances (i.e., non-members) for inference. A widely-used strategy to explore the discrepancy is shadow training [sp17miashokri], which involves training multiple shadow models on datasets drawn from the same distribution as the target model’s training set. For each query instance, the membership scores generated by the shadow models trained without the instance (shadow out models) represent the out distribution for the instance. In some settings, one also has shadow models that are trained with the instance as a member, and can use the scores generated by these shadow in models to compute an in distribution. Given a target model, one tries to tell which distribution the score from the target model fits better. MIAs [iclr23canary, ccs24rapid, ndss26cpmia, icml24rmia, sp22lira] utilizing shadow training have shown strong attack performance.
@@
A critical limitation of existing shadow-based MIAs lies in their substantial computational overhead. Training each shadow model incurs non-trivial overhead, and state-of-the-art attacks like LiRA [sp22lira] and PMIA [ndss26cpmia] require training hundreds of shadow models to estimate the likelihood ratio for inference. This requirement imposes a substantial computational burden, which reduces the feasibility of using these state-of-the-art MIAs for practical privacy auditing [icml24rmia]. In addition, it also impedes research reproducibility within the privacy community (see Section˜2 for further discussion).
We observe that this inefficiency stems from the target-agnostic design of shadow training. That is, the current shadow training process fails to take advantage of knowing the target model under attack. As a result, the shadow models learn only the general patterns of members and non-members, rather than the specific behavioral discrepancies of the target. As shown in the top row of Figure˜1, this target-agnostic approach causes shadow models to exhibit high predictive variance across instances with varying levels of attack vulnerability. Consequently, existing MIAs need to train a multitude of shadow models to capture this variability, resulting in substantial computational overhead and suboptimal performance.
@@
We note that some attacks [ccs22enhanced, tifs25glira, ccs24seqmia] employ model distillation to construct shadow models from the target model. However, these methods directly apply the standard distillation technique [arxiv15distill] without tailoring it to capture the membership-related behaviors of the target model. Furthermore, the resulting distillation models can only reflect behaviors on non-member instances, making it infeasible to exploit behavioral discrepancies for effective membership inference. A more detailed discussion of the differences between our approach and distillation-based methods is provided in Section˜3.2.
We compare the proposed attack with a broad range of MIAs on six benchmark datasets (4 image and 2 non-image datasets), under both non-adaptive and adaptive settings [ndss26cpmia] (also referred to as “offline” and “online” [sp22lira]; see Section˜2 for details). Experimental results show that $\mathsf{IMIA}$ consistently outperforms all evaluated attacks (including distillation-based attacks) while using significantly fewer shadow models, with particularly strong gains in the increasingly advocated non-adaptive setting. For example, $\mathsf{IMIA}$ surpasses the state-of-the-art non-adaptive attack (i.e., PMIA [ndss26cpmia]) by over $\mathbf{9\times}$ in true positive rate at a zero false positive rate on Fashion-MNIST, while using only 5% of its computational resources. Similarly, $\mathsf{IMIA}$ outperforms all adaptive attacks with far less computation. We also conduct comprehensive analyses to show the effectiveness of imitative training over shadow training, evaluate the impact of components in $\mathsf{IMIA}$, and assess the effectiveness of existing defenses against the proposed attack. In summary, we make the following contributions:
@@ 2 Problem Definition and Threat Models
The goal of a membership inference attack (MIA) [sp17miashokri] is to determine whether some instances were included in the training data of a given model $f_{\theta}$. In this paper, we consider the model to be a neural network classifier $f_{\theta}:\mathcal{X}\to\Delta^{c}$, where $f_{\theta}$ is a learned function that maps an input data sample $x\in\mathcal{X}$ to a probability distribution over $c$ classes, where $\Delta^{c}$ denotes the $c$-dimensional simplex. Given a dataset $D$, we denote by $f_{\theta}\leftarrow\mathcal{T}(D)$ the process of training a model $f_{\theta}$, parameterized by weights $\theta$, using a training algorithm $\mathcal{T}$ on dataset $D$.
In early literature [csf18privacy, arixv20revisiting], membership inference is defined via a security game in which the adversary is asked to determine the membership of a single instance. Recent work [ndss26cpmia] refined this definition to more precisely characterize the adversary’s capabilities, particularly whether the adversary is allowed to train shadow models using query instances. Inspired by [ndss26cpmia], we instantiate the membership inference security game to make it match the typical experimental setting of using half members and half non-members for evaluation, and define the following canonical security game:
Definition 1 (Canonical Membership Inference Security Game).
@@
    The challenger randomly selects a member set $D_{\mathrm{a}}\subseteq D$ and samples a nonmember set $D_{\mathrm{b}}$ from $\mathbb{D}$, such that $D_{\mathrm{b}}\cap D=\emptyset$ and $|D_{\mathrm{a}}|=|D_{\mathrm{b}}|$. These two sets are combined to create a query set: $D_{\mathrm{query}}=D_{\mathrm{a}}\cup D_{\mathrm{b}}$, which the challenger then sends to the adversary.
@@
In this paper, we focus on membership inference attacks in black-box scenarios, where the adversary is granted oracle access to the target model $f_{\theta}$, but is not given its parameters. Specifically, the adversary can query the model to obtain softmax output probabilities (using the default softmax temperature of 1) for any input instance. State-of-the-art MIAs [sp22lira, iclr23canary, ccs24rapid, icml24rmia, ndss26cpmia] leverage shadow models [sp17miashokri] to analyze how the model’s outputs change in response to discrepancies between its members versus non-members. We assume that the adversary can construct a dataset by sampling from the data distribution $\mathbb{D}$, and use subsets of this dataset to train shadow models. Depending on when the shadow models are trained, we consider the following two threat models.
@@
Non-Adaptive Setting. In this setting, the adversary is only allowed to train shadow models before the adversary learns the query set $D_{\text{query}}$ (i.e., before step 2 in Definition˜1). That is, the adversary constructs an attack dataset $D_{\text{adv}}^{\text{non-adapt}}$ (from which the adversary will sample subsets to train shadow models) by sampling from $\mathbb{D}$, i.e., $D_{\text{adv}}^{\text{non-adapt}}\sim\mathbb{D}$. For most practical classification tasks, the sizes of $D_{\text{adv}}^{\text{non-adapt}}$ and $D_{\text{query}}$ are relatively small compared to the entire data distribution $\mathbb{D}$, meaning the probability of each instance belonging to both datasets is quite low. As a result, for most query instances, the adversary can only observe the model’s behavior when the query point is not part of the target model’s training set, leading to shadow out models available for inference.
@@
This non-adaptive (offline) setting has attracted growing interest from recent studies [nips23quantile, ccs24seqmia, icml24rmia, ccs24rapid] because it models a realistic attack scenario: the adversary receives a sequence of membership queries and must respond without incurring the cost of retraining shadow models for each query. In this paper, we also follow these studies and mainly focus on the non-adaptive setting.
@@
Adaptive Setting. In this setting, the adversary is allowed to train shadow models after receiving the query set $D_{\text{query}}$ (i.e., after step 2 in Definition˜1). The adversary can therefore leverage $D_{\text{query}}$ for shadow training. For each query instance, the adversary can train both shadow in models (trained with the instance) and shadow out models (trained without it), and exploit their behavioral differences for membership inference.
@@
Recent studies [icml24rmia, nips23quantile, usenix25free] criticize this setting as unrealistic in practice, as it requires training new shadow models for each batch of query instances. Nevertheless, we demonstrate in Section˜3.1 that our proposed method can be adapted to the adaptive setting, achieving state-of-the-art performance while significantly reducing computational overhead.
@@
High Computational Cost in Both Settings. State-of-the-art MIAs in both settings require training a large number of shadow models to achieve their best performance. For instance, LiRA [sp22lira] in the adaptive setting and PMIA [ndss26cpmia] in the non-adaptive setting both suggest training 256 shadow models. The practical implication of this is severe: a single LiRA run on CIFAR-10 can take around 6 days on a modern A100 GPU using its official implementation. Such high computational demands make it infeasible for practical privacy auditing [icml24rmia] and difficult to reproduce research results. It has become common practice for subsequent studies [ccs24seqmia, iclr23canary, ccs24rapid] to use a computationally cheaper version of LiRA (e.g., with 64 or 128 shadow models) as their strongest baseline for evaluation, acknowledging the original version is too costly to run. However, since LiRA’s performance degrades significantly with fewer shadow models in the adaptive setting (as shown in [icml24rmia] and evidenced in Section˜4.2), this compromise leads to skewed evaluations against a weakened baseline.
Assumptions. In this paper, we adopt the standard (albeit sometimes implicit) assumptions used in the MIA literature [ndss26cpmia, sp22lira, iclr23canary, icml24rmia]. Specifically, in the adaptive setting, the adversary is given the membership query set $D_{\text{query}}$, which is constructed using all training instances in $D$ as members and an equal number of non-member instances, i.e., $D_{\mathrm{a}}=D$ and $|D_{\mathrm{b}}|=|D|$ (the equal-size assumption). In the non-adaptive setting, we assume that the adversary can sample a dataset $D_{\text{adv}}^{\text{non-adapt}}$ from the same distribution as the training data, and that $D_{\text{adv}}^{\text{non-adapt}}\cap D_{\text{query}}=\emptyset$ (the disjoint assumption). We emphasize that the equal-size assumption and the disjoint assumption are used only to align with prior experimental setups; the effectiveness of our attack does not depend on them. Discussion of attacks without these assumptions is provided in Appendix B.4.
@@ 3 Imitative Membership Inference Attack
Motivation of Imitative MIA. Our work is motivated by a core limitation in most MIAs: the high predictive variance of membership signals from their target-agnostic shadow models. These shadow models, trained by randomly sampling from the adversary’s dataset, only capture the general patterns of members and non-members across models, failing to account for the specific target model being attacked. To address this, we introduce imitative training, which leverages the target model’s knowledge to train imitative models that explicitly mimic the target’s behavior. This design yields more informative membership signals and, ultimately, more effective attacks. The next subsection describes our approach in detail, and Section˜4.3 demonstrate its effectiveness and efficiency.
@@ 3.1 Attack Method
We begin by introducing the membership signal and the loss function used for imitative training. Next, we outline the imitative training process. Finally, we present the workflow of $\mathsf{IMIA}$, and detail how it can be applied to the adaptive setting.
Scaled Confidence Score. We follow [sp22lira] and define the membership signal of query instance $(x,y)$ on model $f$ as:
@@
where $f(x)_{y}$ denotes the model’s output probability for the true class $y$, and $\max_{y^{\prime}\neq y}f(x)_{y^{\prime}}$ is the highest probability among all incorrect classes. This signal captures the model’s prediction confidence and is widely used in prior MIAs [sp22lira, iclr23canary, ndss26cpmia], as it is a more effective indicator of membership than alternatives such as output loss.
@@
0: Target model $f_{\theta}$, imitation dataset $D_{\text{imitate}}$, pivot dataset $D_{\text{pivot}}$, epochs $T_{\text{warmup}},T_{1},T_{2}$, learning rate $\eta$
@@
5:   if $\text{epoch}\leq T_{\text{warmup}}$ then
6:    # warm-up with cross-entropy loss
7:    $L\leftarrow\frac{1}{|B|}\sum_{(x,y)\in B}\mathcal{L}_{\text{ce}}(x,y,f_{\psi})$
8:   else
9:    $L\leftarrow\frac{1}{|B|}\sum_{(x,y)\in B}\mathcal{L}_{\text{imitate}}(x,y,f_{\psi},f_{\theta})$ $\triangleright$ Eq. (2)
10:   end if
11:   $\psi\leftarrow\psi-\eta\nabla_{\psi}L$
12:  end for
13:  if $\text{epoch}=T_{\text{warmup}}$ then
14:   $f\leftarrow f_{\psi}$ $\triangleright$ warmup checkpoint
15:  end if
@@
18:# continue training to obtain imitative in model for pivots
19: Initialize $f_{\psi}\leftarrow f$
20:for epoch $1,\dots,T_{2}$ do
21:  for each batch $B$ in $D_{\text{pivot}}$ do
22:   $L_{\text{ce}}\leftarrow\frac{1}{|B|}\sum_{(x,y)\in B}\mathcal{L}_{\text{ce}}(x,y,f_{\psi})$ $\triangleright$ cross-entropy loss
23:   $\psi\leftarrow\psi-\eta\nabla_{\psi}L_{\text{ce}}$
24:  end for
25:end for
[…]
27:return $f_{\text{out}},\ f_{\text{in}}^{\text{pivot}}$
@@
Here, the weighting scheme prioritizes the probability corresponding to the ground-truth class, $y$, and the most likely incorrect class, as predicted by the target model. This design encourages imitative models to focus on the most indicative signals, which aligns with our attack signal in Equation˜1.
Our approach can be interpreted as a weighted version of logits matching in model distillation [ijcai21distill], with a weighting strategy specifically tailored to membership inference. In Section˜4.4, we compare it to the commonly used Kullback-Leibler (KL) divergence [arxiv15distill] and also explore other formulations of the imitation loss to demonstrate its effectiveness. A more detailed discussion of the connections to model distillation is provided in the next subsection.
Imitative Training. The imitative training procedure is outlined in Algorithm˜1. It requires query access to the target model and utilizes two datasets selected from the adversary’s data to mimic the target model’s behavior: an imitation dataset $D_{\text{imitate}}$ and a pivot dataset $D_{\text{pivot}}$ (to be detailed later). The training proceeds in two stages. In the first stage, we warm up the model using the standard cross-entropy loss for $T_{\text{warmup}}$ epochs, after which we switch to the proposed imitative loss and train for $T_{1}$ epochs to align the model’s outputs with those of the target model on $D_{\text{imitate}}$ (lines 3–17). Since query instances are not in the adversary’s dataset in the non-adaptive setting, the resulting model serves as an imitative out model, imitating how the target model behaves on non-member inputs. In the second stage, we resume training from the warmup checkpoint for $T_{2}$ epochs on a pre-selected pivot dataset using the standard cross-entropy loss (lines 19–26). This yields an imitative in model for the pivot data, whose behaviors on these instances will serve as proxies to approximate how the target model behaves when a query is in the training set.
@@
The computation of the score $\Lambda$ can be interpreted as a Gaussian likelihood ratio test [sp22lira] with fixed variance, which allows us to frame the inference task as a hypothesis testing. Note that the prepare phase can be done before accessing query instances and only needs to be performed once to answer any queries using the same set of imitative models. During the inference phase, no additional models are trained. The efficiency analysis of each phase is detailed in Section˜4.2.
Selecting Pivot Data. Our attack requires a pivot dataset $D_{\text{pivot}}$ to train imitative in models (line 3 in Algorithm˜2). An effective pivot dataset should (i) cover all classes of the data distribution, (ii) facilitate fast convergence of imitative models to improve efficiency. To construct such a dataset, we adopt a simple yet effective heuristic: we query the target model $f_{\theta}$ with all instances from the adversary’s dataset, compute their losses, and select the $k$ (set as 100 by default) instances with the lowest losses from each class to form $D_{\text{pivot}}$. Our ablation study in Section˜4.4 shows $\mathsf{IMIA}$ is robust to this heuristic. While this simple approach works well in our experiments, more sophisticated pivot-selection strategies may further improve performance; we leave this as future work.
Finding Proxy. During membership inference, we retrieve a set of proxy samples from the pivot dataset (line 18 in Algorithm˜2) to approximate the in behavior of the query instance. Prior work [ndss26cpmia] proposes several proxy-selection heuristics at global, class, and instance granularities. We find that class-level selection consistently offers strong performance while remaining computationally efficient. Therefore, we adopt this strategy: for a given query instance, we retrieve all samples from $D_{\text{pivot}}$ belonging to the same class and use their output distributions to approximate the query’s in behavior.
@@
$\mathsf{IMIA}$ in the Adaptive Setting. While $\mathsf{IMIA}$ is primarily designed for the non-adaptive setting, it can also be applied to the adaptive attack setting. As shown in Algorithm˜3, for a query instance $(x,y)$, we train $N$ pairs of imitative in models (trained with $(x,y)$) and imitative out models (trained without $(x,y)$) on random subsets of the adversary’s dataset $D_{\text{adv}}^{\text{adapt}}$ (lines 2–9). This procedure directly learns the target model’s behavior when the query instance is in its training set, thus eliminating the need for a pivot dataset and proxy selection. The final score is computed by comparing the query’s confidence score on the target model to the mean scores of the imitative in and out distributions (lines 11-14). In addition, while the algorithm is described for a single query, the procedure can be easily parallelized across all query instances using the same set of imitative models. Specifically, we follow [sp22lira] and construct the subsets such that each instance $(x,y)\in D_{\text{query}}$ appears in $N/2$ subsets, and train $N$ imitative models on these subsets. This ensures that the same $N$ imitative models are used to infer the membership for all instances in $D_{\text{query}}$.